# Learning Multiplex Embeddings on Text-rich Networks with One Text Encoder

### v1

@@ Abstract
In real-world scenarios, texts in a network are often linked by multiple semantic relations (e.g., papers in an academic network are referenced by other publications, written by the same author, or published in the same venue). Text documents and their relations form a multiplex text-rich network. Mainstream text representation learning methods use pretrained language models (PLMs) to generate one embedding for each text unit, expecting that all types of relations between texts can be captured by these single-view embeddings. However, this presumption does not hold particularly in multiplex text-rich networks. Along another line of work, multiplex graph neural networks (GNNs) directly initialize node attributes as a feature vector for node representation learning, but they cannot fully capture the semantics of the nodes’ associated texts. In this paper, we propose Metern, a new framework for learning Multiplex Embeddings on TExt-Rich Networks. In contrast to existing methods, Metern uses one text encoder to model the shared knowledge across relations and leverages a small number of parameters per relation to derive relation-specific representations. This allows the encoder to effectively capture the multiplex structures in the network while also preserving parameter efficiency. We conduct experiments on nine downstream tasks in five networks from both academic and e-commerce domains, where Metern outperforms baselines significantly and consistently. Code is available at https://github.com/PeterGriffinJin/METERN-submit.
@@ 1 Introduction
Texts in the real world are often interconnected by multiple types of semantic relations Jin et al. (2023a). For example, papers connected through “same-venue” relation edges lean towards sharing coarse-grained topics, while papers connected through the “cited-by” relation edges tend to share fine-grained topics; E-commerce items linked by “co-viewed” edges usually have related functions, while items linked by “co-brand” edges can have similar designs. The texts and multiple types of links together form a type of network called a multiplex text-rich network, where documents are treated as nodes and the edges reflect multiplex relations among documents. Given a multiplex text-rich network, it is appealing to learn node representations tailored for different relation types, which can be broadly utilized in various downstream tasks (e.g., paper classification and recommendation in academic networks, and item recommendation and price prediction in e-commerce networks) Jin et al. (2023b, c).
@@
A straightforward way to learn node representations in a text-rich network is by encoding the nodes’ associated texts with pretrained language models (PLMs) Brown et al. (2020); Devlin et al. (2019); Liu et al. (2019b). Many studies Cohan et al. (2020); Ostendorff et al. (2022); Reimers and Gurevych (2019) propose to finetune PLMs with contrastive objectives, pushing the text representations with similar semantics to be close in the latent space while pulling those unrelated apart. The semantics correlation between text units is then measured by the similarity of their representations such as with cosine similarity or dot product. However, most existing approaches based on PLMs use a single vector for each text unit, with the implicit underlying assumption that the semantics of different relations between text units are largely analogous, which does not hold universally, particularly in multiplex text-rich networks (demonstrated in Section 3).
@@
To capture the diverse relationships among nodes, multiplex representation learning Jing et al. (2021); Park et al. (2020); Qu et al. (2017); Zhang et al. (2018) is proposed in the graph domain. The philosophy is to learn multiple representations for every node, each encoding the semantics of one relation. They mainly adopt graph neural networks (GNNs) as the backbone model architecture and utilize separate encoders to encode each relation. Nevertheless, these studies with GNNs represent the texts associated with each node as bag-of-words or context-free embeddings Mikolov et al. (2013), which are not enough to characterize the contextualized text semantics. For example, “llama” in biology books and “llama” in machine learning papers should have different meanings given their context, but they correspond to the identical entry in a bag-of-words vector and possess the same context-free embedding. To capture contextualized semantics, a straightforward idea is to adopt PLMs Brown et al. (2020); Devlin et al. (2019); Liu et al. (2019b). However, it is inefficient and unscalable to have separate PLMs for each relation, given that one PLM usually has millions or billions of parameters Kaplan et al. (2020).
@@
To this end, we propose Metern to learn multiplex node/text embeddings with only one shared PLM encoder. The key idea is to introduce “relation prior tokens”, which serve as priors for learning text embeddings. Specifically, relation prior tokens are added upon the original text tokens and fed into a text encoder. The text encoder is encouraged to learn the shared knowledge across different relations, while the relation prior tokens are propelled to capture relation-specific signals. We further explore how to apply Metern to downstream tasks under different scenarios: (1) direct inference where the source relation is clearly singular (e.g., using the “same-author” relation alone for author identification), and (2) indirect inference where source relations might be hybrid and need to be selected via learning. We evaluate Metern on five large-scale networks from the academic domain and the e-commerce domain with nine downstream tasks, where Metern outperforms competitive baseline methods significantly and consistently.
@@
  • Conceptually, we identify the semantic shift across different relations and formulate the problem of multiplex representation learning on text-rich networks.
  • Methodologically, we propose Metern, which learns multiplex text representations with one text encoder and multiple relation prior tokens. Then, we introduce direct inference and “learn-to-select-source-relation” inference with Metern on different downstream tasks.
  • Empirically, we conduct experiments on nine downstream tasks on five datasets from different domains, where Metern outperforms competitive baselines significantly and consistently.
Figure 1: Model Framework Overview. Metern has a language model encoder to model the shared knowledge among relations and relation-prior embeddings to capture the relation-specific signals.
@@ 2 Preliminaries
2.1 Multiplex Text-rich Networks
In a multiplex text-rich network, each node is associated with texts, and nodes are connected by multiple types of edges. We view the texts in each node as a document, and all such documents constitute a corpus $\mathcal{D}$.
Definition 1
(Multiplex Text-rich Networks) A multiplex text-rich network is defined as $\mathcal{G}=\{\mathcal{G}^{1},\mathcal{G}^{2},...,\mathcal{G}^{|\mathcal{R}|}\}=(\mathcal{V},\mathcal{E},\mathcal{D},\mathcal{R})$, where $\mathcal{G}^{r}=(\mathcal{V},\mathcal{E}^{r},\mathcal{D})$ is a network of the relation type $r\in\mathcal{R}$, $\mathcal{V}$ is the set of nodes, $\mathcal{E}=\bigcup_{r\in\mathcal{R}}\mathcal{E}^{r}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of all edges, $\mathcal{D}$ is the set of documents, and $\mathcal{R}$ is the relation type set. Each node $v_{i}\in\mathcal{V}$ is associated with a document $d_{i}\in\mathcal{D}$. Note that $|\mathcal{R}|>1$ for multiplex networks.
@@ 2.2 Problem Definitions
(Learning Multiplex Embeddings on Text-rich Networks) Given a multiplex text-rich network $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{D},\mathcal{R})$, the task of learning multiplex embeddings is to build a model $f_{\Theta}:\mathcal{V}\rightarrow\mathbb{R}^{|\mathcal{R}|\times d}$ with parameters $\Theta$ to learn node representation vectors ${\bm{h}}_{v_{i}}\in\mathbb{R}^{|\mathcal{R}|\times d}$ for each node $v_{i}\in\mathcal{V}$, which should be able to be broadly utilized to various downstream tasks. Note that we aim to learn $|\mathcal{R}|$ embeddings for each node $v_{i}$, with each corresponding to a relation $r\in\mathcal{R}$.
Figure 2: Distribution shift between different relations on Geology network. cb, sa, sv, cr, and ccb represent cited-by, same-author, same-venue, co-reference, and co-cited-by relation respectively. Each entry is the PREC@1 of BERT embedding fine-tuned on the corresponding source relation distribution and tested on the corresponding target relation distribution.
3 Why multiplex embeddings?
@@
Learning a single embedding for each node/document presumes that the embedding ${\bm{h}}_{v}$ is enough to capture the semantics proximity between nodes with a similarity function $\text{Sim}(\cdot)$, i.e., $P(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}},{\bm{h}}_{v_{j}})$, having an underlying assumption that all relations have analogous semantic distributions, i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$ for different relations $r_{k}$ and $r_{l}$. However, this assumption does not always hold in real-world multiplex text-rich networks. For example, two papers ($v_{i},v_{j}$) written by the same author (relation: $r_{k}$) can be either published in the same venue (relation:$r_{l}$) or not; two items ($v_{i},v_{j}$) frequently purchased by the same user (relation: $r_{k}$) can have either the same brand (relation:$r_{l}$) or not.
@@
In addition, we empirically find that the distribution shift across different relations, i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\neq P_{r_{l}}(e_{ij}|v_{i},v_{j})$, in multiplex text-rich networks, truly affects the learned embedding. We finetune BERT [^1] We use the bert-base-uncased checkpoint. Devlin et al. (2019) to generate embeddings on one source relation distribution $P_{r_{k}}(e_{ij}|v_{i},v_{j})$ (row) and test the embeddings on the same or another target relation distribution $P_{r_{l}}(e_{ij}|v_{i},v_{j})$ (column). The results (measured by in-batch PREC@1) are shown in Figure 2. If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly the same, which is not the case in Figure 2. As a result, if we only represent each node/text with one embedding vector, the single representation will mix up the semantics across different relation distributions and will lose accuracy in each of them. This inspires us to learn multiplex embeddings for each node/text, one for each relation, to capture the distinct semantics for each relation. The embeddings belonging to different relations can benefit different downstream tasks, e.g., “same-author” for author identification, and “same-venue” for venue recommendation. More studies on the raw data distributions and learned embeddings can be found in Appendix A.4.
@@ 4 Proposed Method
A straightforward way to learn multiplex embeddings is to have multiple encoders Jing et al. (2021); Park et al. (2020), one for each relation. However, this will make the total number of parameters $|\mathcal{R}|$ times as many as that of a single-encoder model, increasing both memory complexity and time complexity for training and inference, especially when every single encoder is computationally expensive (e.g., pretrained language models Devlin et al. (2019); Liu et al. (2019b)). From another perspective, the semantics of different relations can complement each other, e.g., papers citing each other (i.e., connected via the “cited-by” relation) are highly likely to be of similar topics, thus appearing in the same venue (i.e., also connected via the “same-venue” relation). Nevertheless, adopting separate encoders for each relation will prevent the relations from benefiting each other. To this end, we propose to use only one text encoder to learn multiplex embeddings, with a simple but effective design of “relation prior tokens”. Then, we discuss how to apply the learned embeddings to downstream tasks, including direct inference for tasks where the source relation is clearly singular and “learn to select source relations” for tasks where the source relations might be hybrid. The whole model framework can be found in Figure 1.
4.1 Metern: Learning Multiplex Embeddings with One Text Encoder
Representation Learning with Text Encoder.
Mainstream representation learning models Cohan et al. (2020); Reimers and Gurevych (2019) usually adopt a text encoder $\text{Enc}(\cdot)$ (e.g., Transformer Vaswani et al. (2017)) to generate one node/text embedding for node $v_{i}$, i.e., ${\bm{h}}_{v_{i}}=\text{Enc}(d_{i})$. Then, the predicted similarity score between node $v_{i}$ and $v_{j}$ is calculated by
$P(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}},{\bm{h}}_{v_{j}})=\text{Sim}(\text{Enc}(d_{i}),\text{Enc}(d_{j})),$ (1)
where $\text{Sim}(\cdot,\cdot)$ is a similarity measurement function, e.g., dot product Karpukhin et al. (2020) or cosine similarity Reimers and Gurevych (2019).
However, as we have discussed in Section 3, $P_{r_{k}}(e_{ij}|v_{i},v_{j})\neq P_{r_{l}}(e_{ij}|v_{i},v_{j})$ for different relations $r_{k}$ and $r_{l}$. This motivates us to obtain relation-conditioned embeddings ${\bm{h}}_{v|r}$ rather than just ${\bm{h}}_{v}$ to capture the diverse distribution $P_{r}(e)$ conditioned on relation $r$.

### v2

# Learning Multiplex Representations on Text-Attributed Graphs with One Language Model Encoder

@@ Abstract
In real-world scenarios, texts in a graph are often linked by multiple semantic relations (e.g., papers in an academic graph are referenced by other publications, written by the same author, or published in the same venue), where text documents and their relations form a multiplex text-attributed graph. Mainstream text representation learning methods use pretrained language models (PLMs) to generate one embedding for each text unit, expecting that all types of relations between texts can be captured by these single-view embeddings. However, this presumption does not hold particularly in multiplex text-attributed graphs. Along another line of work, multiplex graph neural networks (GNNs) directly initialize node attributes as a feature vector for node representation learning, but they cannot fully capture the semantics of the nodes’ associated texts. To bridge these gaps, we propose Metag, a new framework for learning Multiplex rEpresentations on Text-Attributed Graphs. In contrast to existing methods, Metag uses one text encoder to model the shared knowledge across relations and leverages a small number of parameters per relation to derive relation-specific representations. This allows the encoder to effectively capture the multiplex structures in the graph while also preserving parameter efficiency. We conduct experiments on nine downstream tasks in five graphs from both academic and e-commerce domains, where Metag outperforms baselines significantly and consistently. Code is available at https://github.com/PeterGriffinJin/METAG.
Machine Learning, ICML
@@ 1 Introduction
Texts in the real world are often interconnected by multiple types of semantic relations. For example, papers connected through the “same-venue” relation edges lean towards sharing coarse-grained topics, while papers connected through the “cited-by” relation edges tend to share fine-grained topics; e-commerce products linked by the “co-viewed” edges usually have related functions, while products linked by the “co-brand” edges can have similar designs. The texts and multiple types of links together form a type of graph called multiplex text-attributed graph (Park et al., 2020), where documents are treated as nodes and the edges reflect multiplex relations among documents. Given such a graph, it is appealing to learn node representations tailored for different relation types, which can be utilized in downstream tasks (e.g., paper classification and recommendation in academic graphs, and item recommendation and price prediction in e-commerce graphs).
@@
A straightforward way to learn node representations in a text-attributed graph is to encode the nodes’ associated texts with pretrained language models (PLMs) (Brown et al., 2020; Devlin et al., 2019; Liu et al., 2019b). Many studies (Cohan et al., 2020; Ostendorff et al., 2022; Reimers & Gurevych, 2019) propose to finetune PLMs with contrastive objectives, pushing the text representations with similar semantics to be close in the latent space while pulling those unrelated apart. The semantics correlation between text units is then measured by the similarity of their representations, such as cosine similarity or dot product. However, most existing approaches based on PLMs use a single vector for each text unit, with the implicit underlying assumption that the semantics of different relations between text units are largely analogous, which does not hold universally, particularly in multiplex text-attributed graphs (shown in Section 3).
@@
To capture the diverse relationships among nodes, multiplex representation learning (Jing et al., 2021; Park et al., 2020; Qu et al., 2017; Zhang et al., 2018) is proposed in the graph domain. The philosophy is to learn multiple representations for every node, each encoding the semantics of one relation. They mainly adopt graph neural networks (GNNs) as the backbone model architecture and utilize separate encoders to encode each relation. Nevertheless, these studies with GNNs represent the texts associated with each node as bag-of-words or context-free embeddings (Mikolov et al., 2013), which are not enough to characterize the contextualized text semantics. For example, “llama” in biology books and “llama” in machine learning papers should have different meanings given their context, but they correspond to the identical entry in a bag-of-words vector and possess the same context-free embedding. To capture contextualized semantics, a straightforward idea is to adopt PLMs (Brown et al., 2020; Devlin et al., 2019; Liu et al., 2019b). However, it is inefficient and unscalable to have separate PLMs for each relation, given that one PLM usually has millions or billions of parameters (Kaplan et al., 2020).
@@
To this end, we propose Metag to learn multiplex node/text embeddings with only one shared PLM encoder. The key idea is to introduce “relation prior tokens”, which serve as priors for learning text embeddings. Specifically, relation prior tokens are prepended to the original text tokens and fed into a text encoder. The text encoder is encouraged to learn the shared knowledge across different relations, while the relation prior tokens are propelled to capture relation-specific signals. We further explore how to apply Metag to downstream tasks under different scenarios: (1) direct inference where the source relation is clearly singular (e.g., using the “same-author” relation alone for author identification), and (2) indirect inference where source relations might be hybrid and need to be selected via learning. We evaluate Metag on five large-scale graphs from the academic domain and the e-commerce domain with nine downstream tasks, where Metag outperforms competitive baseline methods significantly and consistently.
@@
  • Conceptually, we identify the semantic shift across different relations and formulate the problem of multiplex representation learning on text-attributed graphs.
  • Methodologically, we propose Metag, which learns multiplex text representations with one text encoder and multiple relation prior tokens. Then, we introduce direct inference and “learn-to-select-source-relation” inference with Metag on different downstream tasks.
  • Empirically, we conduct experiments on nine downstream tasks on five datasets from different domains, where Metag outperforms competitive baselines significantly and consistently.
@@ 2 Preliminaries
2.1 Multiplex Text-Attributed Graphs
In a multiplex text-attributed graph, each node is associated with texts, and nodes are connected by multiple types of edges. We view the texts in each node as a document, and all such documents constitute a corpus $\mathcal{D}$.
Definition 2.1.
(Multiplex Text-Attributed Graphs) A multiplex text-attributed graph is defined as $\mathcal{G}=\{\mathcal{G}^{1},\mathcal{G}^{2},...,\mathcal{G}^{|\mathcal{R}|}\}=(\mathcal{V},\mathcal{E},\mathcal{D},\mathcal{R})$, where $\mathcal{G}^{r}=(\mathcal{V},\mathcal{E}^{r},\mathcal{D})$ is a graph of the relation type $r\in\mathcal{R}$, $\mathcal{V}$ is the set of nodes, $\mathcal{E}=\bigcup_{r\in\mathcal{R}}\mathcal{E}^{r}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of all edges, $\mathcal{D}$ is the set of documents, and $\mathcal{R}$ is the relation type set. Each node $v_{i}\in\mathcal{V}$ is associated with a document $d_{i}\in\mathcal{D}$. Note that $|\mathcal{R}|>1$ for multiplex graphs (Park et al., 2020).
@@ 2.2 Problem Definitions
(Learning Multiplex Representations on Text-Attributed Graphs) Given a multiplex text-attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{D},\mathcal{R})$, the task of learning multiplex embeddings is to build a model $f_{\Theta}:\mathcal{V}\rightarrow\mathbb{R}^{|\mathcal{R}|\times d}$ with parameters $\Theta$ to learn node representation vectors ${\bm{h}}_{v_{i}}\in\mathbb{R}^{|\mathcal{R}|\times d}$ for each node $v_{i}\in\mathcal{V}$, which should be able to be broadly utilized to various downstream tasks. Note that we aim to learn $|\mathcal{R}|$ embeddings for each node $v_{i}$, with each corresponding to a relation $r\in\mathcal{R}$.
3 Why multiplex representations?
@@
Learning a single embedding for each node/document presumes that the embedding ${\bm{h}}_{v}$ is enough to capture the semantics proximity between nodes with a similarity function $\text{Sim}(\cdot)$, i.e., $P(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}},{\bm{h}}_{v_{j}})$, having an underlying assumption that all relations have analogous semantic distributions, i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$ for different relations $r_{k}$ and $r_{l}$. However, this assumption does not always hold in real-world multiplex text-attributed graphs. For example, two papers ($v_{i},v_{j}$) written by the same author (relation: $r_{k}$) can be either published in the same venue (relation:$r_{l}$) or not. We measure the raw data semantic distribution shift between relations with the Jaccard score:
$\text{Jac}(r_{k},r_{l})=\frac{|P_{r_{k}}(e_{ij})\cap P_{r_{l}}(e_{ij})|}{|P_{r_{k}}(e_{ij})\cup P_{r_{l}}(e_{ij})|}.$ (1)
(a) Raw Data
(b) Performance
Figure 1: (a) Raw data distribution shift and (b) performance shift across different relations on a graph of Geology papers. cb, sa, sv, cr, and ccb represent cited-by, same-author, same-venue, co-reference, and co-cited-by relation, respectively. Each entry in (a) and (b) are the Jaccard score and the PREC@1 of BERT embedding fine-tuned on the corresponding source relation distribution and tested on the corresponding target relation distribution.
@@
The results on the graph of Geology papers are shown in Figure 1(a). If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly 1, which is not the case in Figure 1(a). In addition, we empirically find that the distribution shift across different relations, i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\neq P_{r_{l}}(e_{ij}|v_{i},v_{j})$, in multiplex text-attributed graphs truly affects the learned embedding. We finetune BERT [^1] We use the bert-base-uncased checkpoint. (Devlin et al., 2019) to generate embeddings on one source relation distribution $P_{r_{k}}(e_{ij}|v_{i},v_{j})$ (row) and test the embeddings on the same or another target relation distribution $P_{r_{l}}(e_{ij}|v_{i},v_{j})$ (column). The results (measured by in-batch PREC@1) are shown in Figure 1(b). If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be almost identical, which is not the case in Figure 1(b). As a result, if we only represent each node/text with one embedding vector, the single representation will mix up the semantics across different relation distributions and will lose accuracy in each of them. This inspires us to learn multiplex embeddings for each node/text, one for each relation, to capture the distinct semantics for each relation. The embeddings belonging to different relations can benefit different downstream tasks, e.g., “same-author” for author identification, and “same-venue” for venue recommendation. More studies on the raw data distributions and learned embeddings can be found in Appendix A.4.
@@ 4 Proposed Method
Figure 2: Model Framework Overview. Metag has a language model encoder to model the shared knowledge among relations and relation-prior embeddings to capture the relation-specific signals.
A straightforward way to learn multiplex embeddings is to have multiple encoders (Jing et al., 2021; Park et al., 2020), one for each relation. However, this will make the total number of parameters $|\mathcal{R}|$ times as many as that of a single-encoder model, increasing both memory complexity and time complexity for training and inference, especially when every single encoder is computationally expensive (e.g., PLMs (Devlin et al., 2019; Liu et al., 2019b)). From another perspective, the semantics of different relations can complement each other, e.g., papers citing each other (i.e., connected via the “cited-by” relation) are highly likely to be of similar topics, thus appearing in the same venue (i.e., also connected via the “same-venue” relation). Nevertheless, adopting separate encoders for each relation will prevent the relations from benefiting each other. To this end, we propose to use only one language model encoder to learn multiplex embeddings, with a simple but effective design of “relation prior tokens”. We also propose flexible ways to apply the learned embeddings to downstream tasks, including direct inference for tasks where the source relation is clearly singular and “learn to select source relations” for tasks where the source relations might be hybrid. The whole model framework can be found in Figure 2.
4.1 Metag: Learning Multiplex Embeddings with One Language Model Encoder
Representation Learning with Language Model Encoder.
Mainstream representation learning models (Cohan et al., 2020; Reimers & Gurevych, 2019) usually adopt a language model encoder $\text{Enc}(\cdot)$ (e.g., Transformer (Vaswani et al., 2017)) to generate one node/text embedding for node $v_{i}$, i.e., ${\bm{h}}_{v_{i}}=\text{Enc}(d_{i})$. Then, the predicted similarity score between node $v_{i}$ and $v_{j}$ is calculated by
$\small P(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}},{\bm{h}}_{v_{j}})=\text{Sim}(\text{Enc}(d_{i}),\text{Enc}(d_{j})),$ (2)
where $\text{Sim}(\cdot,\cdot)$ is a similarity measurement function, e.g., dot product (Karpukhin et al., 2020) or cosine similarity (Reimers & Gurevych, 2019).
However, as discussed in Section 3, $P_{r_{k}}(e_{ij}|v_{i},v_{j})\neq P_{r_{l}}(e_{ij}|v_{i},v_{j})$ for different relations $r_{k}$ and $r_{l}$. This motivates us to obtain relation-conditioned embeddings ${\bm{h}}_{v|r}$ rather than just ${\bm{h}}_{v}$ to capture the diverse distribution $P_{r}(e)$ conditioned on relation $r$.